Four names are put on a ballot in randomly select order. What is the probability that they are NOT in alphabetical order?
Answer by jim_thompson5910(35256)

There are 4! = 4*3*2*1 = 24 ways to order the names. Only one way is exactly in alphabetical order. The other 24-1 = 23 ways are NOT in alphabetical order (whether it's one name out of place or all of them are)

So the probability is 23/24